16. (16 points) Suppose the displacement u(x, t) of a piece of flexible string is given by the initial- boundary value problePlease just explain parts d and f. Part d is 0 and f is true
Answer #1

function of ne Wheree 100 Unexe = Utt — 100 et les 100 meteen Trial solution is u = xT --- ® x is a - T is a function of t 24- m2 x = -mx x = 0,02 + 2 Si2x I = -m2 the-m7.100 = c cos lomt te simtomt) Som is UCX,T)= X.T u = C, coome+Casinma) .( Gcool22 (46=0 22 u(4,4)=0 x240 U=0, x=4 Os (2 sin 4m ſ 6 consomt +44 sintomt] sin 4m-0 Am ФЛ 1 Where na 1, 2, 3 ... 4 Turx.7) = C,TIA 2) u(x,0)=0 t=0, U=0 Os C2 sinmäe [ 33+] sinma = 0 ma=nT Im=1 cs Scanned with CamScanner